OpenStudy (anonymous):

Help please. B-(2b-7)=-3

OpenStudy (shamil98):

b -(2b-7) = -3 b -2b + 7 = -3 -b = -10 b = 10

OpenStudy (shamil98):

b - (2b-7) you distribute the negative 1 to 2b and -7 which is why it is -2b and +7

OpenStudy (shamil98):

From there you just solve for b .

OpenStudy (anonymous):

Oh. I see! sorry its my first time learning this!